OpenStudy (anonymous):

write the equation in slope -intercept form that contains the point (4,-7) and (0,5)

OpenStudy (anonymous):

first find the slope... How?

OpenStudy (anonymous):

(y2-y1) -------- = m (x2-x1)

OpenStudy (anonymous):

k...slop is -3

OpenStudy (anonymous):

slope intercept is y = -3x-5

OpenStudy (anonymous):

y2 - y1 / x2- x1 is how you find slope so it's 5+7/ 0-4 = 12/-4 = -3

OpenStudy (anonymous):

are you sure that's the intercept form? :)

OpenStudy (anonymous):

you need to use point slope formula after you find slope...

OpenStudy (anonymous):

then use the point slope formula to get it in slpe intercept form.... y - y1 = m (x-x1) y +7 = -3 (x-4) y+ 7 = -3x + 12 y = -3x +5

OpenStudy (anonymous):

thats the wright answer :) thank you

OpenStudy (anonymous):

your welcome